# On the Parallelizability of Approval-Based Committee Rules

### v1

@@ C.4 Method of Equal Shares
Suppose we are given candidates $C$, collection of votes $V$, and committee size $k$. MES+seq-P runs in two phases. In the first phase, we run MES: we select $k^{\prime}$ candidates as described in Definition 9. If $k^{\prime}<k$, then we run seq-Phragmén on the remaining candidates to pick a committee of size $k^{\prime}-k$, where the loads of each voter $i$ are initialized by setting $y_{0}(i)=-x_{k^{\prime}}(i)$.
@@
Suppose we are given an instance $(G,v)$ of LFMIS, where $G$ is a 3-regular graph with $m$ vertices labeled $1$ to $m$. We will construct an election with candidate $C$, collection of votes $V$, and committee size $k$, and specify a candidate $c\in C$ such that the winning committee selected by MES+seq-P includes $c$ if and only if $v$ is included in the LFMIS of $G$.
As in our previous proof, we add to $C$ each vertex in $G$ and add to $V$ a vote for each edge in $G$. We will also add many copies of candidates and votes equivalent to a 13-vertex tree graph, denoted $T$: $T$ has a root vertex with four children, each of which has two children. We add $13m$ “copies of $T$” to our election system, i.e., we add $13m$ extra candidates to $C$, labeled $m+1$ to $14m$, and $12m$ votes to $V$ equivalent to the edges of $T$. Let $k=m/2+4m$ and $c=v$.
@@
Recall that initially we are running MES. We have $n=3m/2+12m$ votes, so the initial budget for each voter is $k/n=(m/2+4m)/(3m/2+12m)=1/3$. Observe that the candidates corresponding to the root of $T$ have four voters each. So, initially, all of these roots can be elected by incurring a cost of $1/4$ from each of their voters. Since this is the minimum cost, $\rho_{c}$, MES will pick each root for the winning committee, leaving each root’s voters with budget $1/3-1/4=1/12$. After $m$ roots have been elected, observe that no other candidate corresponding to a vertex from $T$ can be elected by MES; those in the second layer (children of the roots) have voters with budget $1/3+1/3+1/12<1$, and those in the last layer (leaf vertices of $T$) have single voters with budget $1/3$. We are now in a situation similar to the one in the proof of Theorem 11; the candidates corresponding to vertices in $V(G)$ can be elected by incurring a cost of $1/3$ on each of their voters. As argued in said proof, the candidates corresponding exactly to the LFMIS of $V(G)$ will be elected. Once all of these candidates have been elected, no candidates’ voters have the budget to elect them and the first phase concludes.
@@
$\ell_{1}(c)=\frac{1+\left(-\frac{1}{3}-\frac{1}{3}\right)}{3}=\frac{1}{9}\text{ or }\ell_{1}(c)=\frac{1+\left(-\frac{1}{3}-\frac{1}{3}\right)}{3}=\frac{2}{9}$
@@
$\ell_{1}(c)=\frac{1+\left(-\frac{1}{3}\right)}{3}=\frac{2}{3}$
@@
Note that alternatively one could easily reduce seq-Phragmén to MES+seq-P by adding enough dummy votes so that the first phase is skipped. However that result alone does not give us insight into the hardness of MES.

### v3

@@ C.4 Method of Equal Shares + Phragmén
Suppose we are given candidates $C$, collection of votes $V$, and committee size $k$. MES+seq-P runs in two phases. In the first phase, we run MES: we select $k^{\prime}$ candidates as described in Definition 10. If $k^{\prime}<k$, then we run seq-Phragmén on the remaining candidates to pick a committee of size $k-k^{\prime}$, where the loads of each voter $i$ are initialized by setting $y_{0}(i)=-x_{k^{\prime}}(i)$.
@@
Suppose we are given an instance $(G,v)$ of LFMIS, where $G$ is a 3-regular graph with $m$ vertices labeled $1$ to $m$. We will construct an election with candidates $C$, collection of votes $V$, and committee size $k$, and specify a candidate $c\in C$ such that the winning committee selected by MES+seq-P includes $c$ if and only if $v$ is included in the LFMIS of $G$.
As in our previous proof, we add to $C$ each vertex in $G$ and add to $V$ a vote for each edge in $G$. We will also add many copies of candidates and votes equivalent to a 13-vertex tree graph, denoted $T$: $T$ has a root vertex with four children, each of which has two children. We add $13m$ “copies of TT” to our election system, i.e., we add $13m$ extra candidates to $C$, labeled $m+1$ to $14m$, and $12m$ votes to $V$ equivalent to the edges of $T$. Let $k=m/2+4m$ and $c=v$.
@@
Recall that initially we are running MES. We have $n=3m/2+12m$ votes, so the initial budget for each voter is $k/n=(m/2+4m)/(3m/2+12m)=1/3$. Observe that the candidates corresponding to the root of $T$ have four voters each. So, initially, all of these roots can be elected by incurring a cost of $1/4$ from each of their voters. Since this is the minimum cost, $\rho_{c}$, MES will pick each root for the winning committee, leaving each root’s voters with budget 1/31/4=1/121/3-1/4=1/12. After $m$ roots have been elected, observe that no other candidate corresponding to a vertex from $T$ can be elected by MES; those in the second layer (children of the roots) have voters with budget $1/3+1/3+1/12<1$, and those in the last layer (leaf vertices of $T$) have single voters with budget $1/3$. We are now in a situation similar to the one in the proof of Theorem 12; the candidates corresponding to vertices in $V(G)$ can be elected by incurring a cost of $1/3$ on each of their voters. As argued in said proof, the candidates corresponding exactly to the LFMIS of $V(G)$ will be elected. Once all of these candidates have been elected, no candidates’ voters have the budget to elect them and the first phase concludes.
@@
$\ell_{1}(c)=\frac{1+\left(-\frac{1}{3}-\frac{1}{3}\right)}{3}=\frac{1}{9}\text{ or }\ell_{1}(c)=\frac{1+\left(-\frac{1}{3}\right)}{3}=\frac{2}{9}$
@@
$\ell_{1}(c)=\frac{1+\left(-\frac{1}{3}\right)}{3}=\frac{2}{9}$
@@
C.5 MES without tie-breaking
Theorem 13. Computing the winning committee for (the first phase of) MES is inherently sequential even when tie-breaking is never invoked.
Proof.
Suppose we are given an instance $(G,v)$ of LFMIS, where $G$ is a 3-regular graph with $m$ vertices labeled $1$ to $m$. Without loss of generality, we will assume that $m\geq 12$. We will construct an election with candidate $C$, collection of votes $V$, and committee size $k$, and specify a candidate $c\in C$ such that the winning committee selected by MES includes $c$ if and only if $v$ is included in the LFMIS of $G$.
Let $C$ have a candidate for each vertex of $G$. For each $\{a,b\}\in E(G)$, we add $m^{2}$ votes $\{a,b\}$ to $V$. We also add (ma)(m-a) votes for each $a\in C$. Note that each candidate $a$ has $3m^{2}+(m-a)$ votes. Finally, add an extra candidate, labeled 0, that has $m^{3}$ singleton votes. It is clear that there are no ties in the election. Note that we have a total of $n=3m^{3}/2+m(m+1)/2+m^{3}$ votes. Finally, we set $c=v$ and $k=m+1$. Observe that the initial budget is
$x_{0}(i)=\frac{k}{n}=\frac{m+1}{3m^{3}/2+m(m-1)/2+m^{3}}$
for each voter $i$. We now show that the candidates corresponding to $V(G)$ picked by MES correspond exactly to the vertices in the LFMIS of $G$. We prove this statement via the following two claims for all candidates other than candidate 0.
Claim 36.
For an unelected candidate $a>0$, if its voters have not already voted for another candidate, $a$’s voters have the budget to elect $a$.
Proof.
A candidate can be elected in round $r$ if $\sum_{i\in N(a)}x_{r}(i)\geq 1$. If $a$’s voters have not voted for anyone else, then $x_{r}(i)=x_{0}(i)$. Thus, for each candidate $a$, we have
$\sum_{i\in N(a)}x_{0}(i)$ $=x_{0}(i)\times(3m^{2}+(m-a))$
$\geq x_{0}(m)\times(3m^{2}+(m-m))$
$\geq\frac{3m^{2}(m+1)}{3m^{3}/2+m(m-1)/2+m^{3}}$
The R.H.S. is $\geq 1$ for all $m\geq 1$. ∎
Claim 37.
For candidates $a>0$ and $b>0$ that share a vote, if $b$ was selected by the MES in a previous round, then the voters of $a$ do not have the budget to elect $a$.
Proof.
Candidate $b$ has $3m^{2}+(m-b)$ votes. The minimum amount that each voter spent to elect $b$ is $1/(3m^{2}+(m-b))$. Of these votes, $m^{2}$ are shared with $a$. Thus, the maximum amount of total budget that the voters of $a$ have is
$\sum_{i\in N(a)}x_{r}(i)$ $\leq(2m^{2}+(m-a))x_{0}(i)+\left(x_{0}(i)-\frac{1}{(3m^{2}+(m-b))}\right)m^{2}$
$\leq(2m^{2}+m)x_{0}(i)+\left(x_{0}(i)-\frac{1}{(3m^{2}+m)}\right)m^{2}$
Substituting $x_{0}(i)$ in the R.H.S. and simplifying gives us the following expression:
$\frac{2(17m+8)}{5(5m^{2}+m-1)}+\frac{1}{3(3m+1)}+\frac{13}{15}$
This is clearly decreasing as $m$ grows, and it can be verified to be <1<1 for all $m\geq 12$. ∎
Recall that MES prefers picking the candidate that minimizes the maximum cost incurred by each voter. Candidate 0 is picked first since it has the most votes and the cost incurred is split among these voters. Similar reasoning shows that, candidate 1 is picked since it has the most voters ($3m^{2}+m-1$). By Claims 36 and 37 and the fact that candidates with more votes (i.e., vertices with earlier lexicographic order) are preferred, it is clear that the candidates picked after candidate 1 will correspond exactly to the vertices in the LFMIS of $G$. Since we can pick up to m1m-1 remaining candidates, we are guaranteed to pick all the candidates corresponding to the LFMIS of $G$.
Note that the extra padding of $m^{3}$ votes given by candidate 0 is what allows us to be in the “goldilocks zone” where the budget is large enough so that every candidate has large enough initial budget to be elected, but taking away some of the budget of $m^{2}$ of its votes does not allow it to be elected.